Does the negative exponent rule for math come into play for chemistry?

When dealing with unit factors and conversions

I want to put cm and m on a unit factor, do I just make the exponent positive.


m/10^2cm or 10^2/m
Thank you
Answer #1

The negative exponents rule is applicable wherever there is any calculation, and since a lot of calculations are there in chemistry as well, so yes you have to apply it whenever you see a chance.

1 m = 100 cm

And,

1 cm = 10-2 m = (1/100) m

Hope this helps !